# Effective models for dense vortex lattices in the Kitaev honeycomb model

### 22 Dec 2023 v1

@@ Appendix A Validity of the effective models
At small $\kappa$, the energy splitting of two isolated MZMs separated by a distance $d$ is qualitatively given by $t(d,\kappa)\simeq\Delta_{\rm b}e^{-d/\xi}\cos(k_{\rm F}d+\gamma)$, where $d=|\bm{A}_{1}|=|\bm{A}_{2}|$ is the vortex separation, $\xi$ is the superconducting coherence length ($\xi\sim W_{\rm b}/\Delta_{\rm b}$ with $W_{\rm b}$ the total bandwidth of the background), $k_{F}$ is the distance from the center of the Brillouin zone to the Dirac points and $\gamma$ is a phaseshift [30, 12]. These Friedel-like oscillations are a hallmark of the two-vortex solution of isolated MZMs, and naturally appear in other systems exhibiting Ising anyons, such as the Moore-Read quantum Hall state, $p_{x}+ip_{y}$ superconductors, and topological nanowires. The oscillations are related to the change in fermionic parity and to the fact that MZMs behave as Ising anyons $\sigma$ that have the possibility to fuse either in the vacuum $1$ or in a fermion $\psi$ according to the well-known fusion rule $\sigma\times\sigma=1+\psi$ (see Fig. 1 in Ref. [12]). We may estimate the total vortex bandwidth $\delta_{\rm v}$ for the direct vortex configurations by $\delta_{\rm v}\sim|t(d,\kappa)|\sim\Delta_{\rm b}e^{-d/\xi}$. Dilute vortex configurations have MZMs well-separated from one another, i.e. $d\gg\xi$. This implies that $\delta_{\rm v}\ll\Delta_{\rm b}$, which we take as a necessary condition for the validity of the effective model for the direct configurations.
@@
For the dual vortex configurations, the low-energy vortex bandwidth $\delta_{\rm v}$ may be estimated as the energy level splitting $2\epsilon$ of the two Majorana modes bound to an isolated white plaquette in a vortex-full sea. Figure 6 of Ref. [15] shows $2\epsilon$ to always be less than the vortex-full background’s energy gap $\Delta_{\rm b}$, except near $\kappa=0$ and $\kappa=1/2$ where $2\epsilon$ and $\Delta_{\rm b}$ vanish. Thus, the $\delta_{\rm v}\ll\Delta_{\rm b}$ condition is essentially equivalent to guaranteeing that $\kappa\neq 0$ and $\kappa\neq 1/2$. Dual vortex configurations that are “dilute” actually mean that the vortex density $\rho=(n-1)/n\to 1$.
In Fig. 11, we plot the ratio $\delta_{\rm v}/\Delta_{\rm b}$ for the direct $\rho=1/21$ and dual odd $\rho=20/21$ vortex configurations as a function of $\kappa$. We offer these two cases as representative examples, as other choices of direct and dual configuration densities exhibit the same qualitative features. As a visual guide, we draw a dotted line at $\delta_{\rm v}/\Delta_{\rm b}=1/2$ to indicate the regions where $\delta_{\rm v}/\Delta_{\rm b}\ll 1$. For the direct configuration in Fig. 11(a), we see that the ratio $\delta_{\rm v}/\Delta_{\rm b}$ is equal to $1/2$ for the two critical values $\kappa_{c_{1}}$ and $\kappa_{c_{2}}$, represented by the red circle and triangle, respectively. In Table 1 we see that the energy gap of the vortex-free ($\rho=0$) background (roughly equal to $\Delta_{\rm b}$) vanishes as $\kappa=0$ and is of order $1$ as $\kappa\to\infty$. Thus, the low-energy vortex bands are only well-separated from the background as $\kappa_{c_{1}}<\kappa<\kappa_{c_{2}}$. For less dense triangular vortex configurations in the $\rho\to 0$ limit, we find that this region of validity increases, such that $\kappa_{c_{1}}\to 0$ and $\kappa_{c_{2}}\to\infty$.
@@ B.1 One contact point
This type of transition occurs when the vortex Chern number changes from $\nu_{\rm v}=-1$ to $\nu_{\rm v}=0$ for the dual odd vortex configurations of density $\rho=(n-1)/n$ such that $n$ is not a multiple of three. That is, for densities such as $\rho=12/13$, $18/19$, $24/25$, and etc. These transitions generally occur when $\kappa_{c}\sim 0.1-0.2$. Before the transition ($\kappa<\kappa_{c}$), the system features two gapped Dirac points of opposite chirality and is a Chern insulator with $\nu_{\rm v}=-1$. As $\kappa$ increases, the Dirac points move and at a critical $\kappa_{c}$ they meet at a time-reversal invariant momentum (TRIM) where they annihilate and the gap closes. Exactly at the transition, there is a single contact point that does not carry a topological charge. After this merging transition, the gap reopens and the system becomes a trivial insulator with $\nu_{\rm v}=0$. This merging transition is similar to the one occurring in the Kitaev honeycomb model when evolving from the isotropic point $J_{x}=J_{y}=J_{z}$ to a very anisotropic case [5], or also in uniaxially-strained graphene [31].
@@
We note that this Hamiltonian results from a different gauge choice for the $s_{jk}^{(\alpha)}$, $s_{jk}^{(\perp)}$, and $\widetilde{s}_{jk}^{(\perp)}$ variables than what is shown in Fig. 8. In Fig. 12 we plot the dispersion of both this effective model and the low-lying vortex bands of the Kitaev honeycomb model of the $\nu_{\rm v}=-1\to 0$ transition in the $\rho=12/13$ dual odd vortex configuration as $\kappa_{c}=0.219461$ [15]. The proper parameter values for the effective model are given in the caption. The critical point of the effective model appears when $t_{\perp 1}=\epsilon/6$. As $t_{\perp 1}<\epsilon/6$ the Chern number of the effective model is $\nu_{\rm v}=-1$, and when $t_{\perp 1}>\epsilon/6$ the Chern number is $\nu_{\rm v}=0$. The sign of $t_{\perp 2}$ controls the sign of $\nu_{\rm v}$, and the choice of signs of the cosine terms of $\widetilde{g}_{1}(\bm{k})$ controls the position of the contact point within the Brillouin zone. If we were to instead analyze a different vortex density exhibiting a one contact point transition, one may have to alter the signs of the cosine terms in order to properly match the location of the contact point within the Brillouin zone (this is true for the two and six contact points models as well). Note that because the value of $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.38$ is relatively high (i.e. the vortex bands are not so well separated from the background bands), the bandwidth of the effective model overestimates that of the low-lying vortex bands of the Kitaev honeycomb model.
@@ B.2 Two contact points
In Fig. 13 we plot the dispersion of both this effective model and the low-lying vortex bands of the Kitaev honeycomb model of the $\nu_{\rm v}=-2\to 0$ transition in the $\rho=20/21$ dual odd vortex configuration, as $\kappa_{c}=0.106486$. The proper parameter values for the effective model are given in the caption. The critical points of the effective model appear when $t_{\perp 1}=-\epsilon/3$. As $t_{\perp 1}>-\epsilon/3$ the Chern number of the effective model is $\nu_{\rm v}=0$, and when $t_{\perp 1}<-\epsilon/3$ the Chern number is $\nu_{\rm v}=-2$. The sign of $t_{2}$ controls the sign of $\nu_{\rm v}$. Note that $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.62$, and hence the bandwidth of the effective model is slightly larger than that of the low-lying vortex bands of the Kitaev honeycomb model.
We note that in this system, the energy spectrum is gapped and topologically nontrivial only if the sign of one of the sine terms in $\widetilde{g}_{2}(\bm{k})$ is dissimilar to the sign of the other two. If every sine term were positive, for example, then the $2\times 2$ matrix Hamiltonian may be expressed as the linear combination of only two Pauli matrices and would either be gapless before the transition (i.e. a topological semi-metal with gapless Dirac points) or gapped after the transition with a trivial Chern number of $\nu_{\rm v}=0$, due to a merging transition.
@@ B.3 Six contact points
In Fig. 14 we plot the dispersion of both this effective model and the low-lying vortex bands of the Kitaev honeycomb model of the $\nu_{\rm v}=-3\to 3$ transition in the $\rho=6/7$ dual odd vortex configuration, as $\kappa_{c}=1.80057$. The proper parameter values for the effective model are given in the caption. The critical points of the effective model appear when $t_{\perp 2}=0$. As $t_{\perp 2}>0$ the Chern number of the effective model is $\nu_{\rm v}=-3$, and when $t_{\perp 2}<0$ the Chern number is $\nu_{\rm v}=3$. The ratio $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.38$ indicates that the bandwidth of the effective model is slightly larger than that of the low-lying vortex bands of the Kitaev honeycomb model. Note that because $t_{1}\neq t_{2}=0$, we do not have the same issue of the sine’s signs as discussed in the two contact points section, and the $2\times 2$ matrix Hamiltonian is still composed of three different Pauli matrices.

### 05 Mar 2024 v2

@@ Appendix A Validity of the effective models
At small $\kappa$, the energy splitting of two isolated MZMs separated by a distance $d$ is qualitatively given by $t(d,\kappa)\simeq\Delta_{\rm b}e^{-d/\xi}\cos(k_{\rm F}d+\gamma)$, where $d=|\bm{A}_{1}|=|\bm{A}_{2}|$ is the vortex separation, $\xi$ is the superconducting coherence length ($\xi\sim W_{\rm b}/\Delta_{\rm b}$ with $W_{\rm b}$ the total bandwidth of the background), $k_{F}$ is the distance from the center of the Brillouin zone to the Dirac points and $\gamma$ is a phaseshift [33, 12]. These Friedel-like oscillations are a hallmark of the two-vortex solution of isolated MZMs, and naturally appear in other systems exhibiting Ising anyons, such as the Moore-Read quantum Hall state, $p_{x}+ip_{y}$ superconductors, and topological nanowires. The oscillations are related to the change in fermionic parity and to the fact that MZMs behave as Ising anyons $\sigma$ that have the possibility to fuse either in the vacuum $1$ or in a fermion $\psi$ according to the well-known fusion rule $\sigma\times\sigma=1+\psi$ (see Fig. 1 in Ref. [12]). We may estimate the total vortex bandwidth $\delta_{\rm v}$ for the direct vortex configurations by $\delta_{\rm v}\sim|t(d,\kappa)|\sim\Delta_{\rm b}e^{-d/\xi}$. Dilute vortex configurations have MZMs well-separated from one another, i.e., $d\gg\xi$. This implies that $\delta_{\rm v}\ll\Delta_{\rm b}$, which we take as a necessary condition for the validity of the effective model for the direct configurations.
@@
For the dual vortex configurations, the low-energy vortex bandwidth $\delta_{\rm v}$ may be estimated as the energy level splitting $2\epsilon$ of the two Majorana modes bound to an isolated white plaquette in a vortex-full sea. Figure 6 of Ref. [17] shows $2\epsilon$ to always be less than the vortex-full background’s energy gap $\Delta_{\rm b}$, except near $\kappa=0$ and $\kappa=1/2$, where $2\epsilon$ and $\Delta_{\rm b}$ vanish. Thus the $\delta_{\rm v}\ll\Delta_{\rm b}$ condition is essentially equivalent to guaranteeing that $\kappa\neq 0$ and $\kappa\neq 1/2$. Dual vortex configurations that are “dilute” actually mean that the vortex density $\rho=(n-1)/n\to 1$.
In Fig. 11, we plot the ratio $\delta_{\rm v}/\Delta_{\rm b}$ for the direct $\rho=1/21$ and dual odd $\rho=20/21$ vortex configurations as a function of $\kappa$. We offer these two cases as representative examples, as other choices of direct and dual configuration densities exhibit the same qualitative features. As a visual guide, we draw a dotted line at $\delta_{\rm v}/\Delta_{\rm b}=1/2$ to indicate the regions where $\delta_{\rm v}/\Delta_{\rm b}\ll 1$. For the direct configuration in Fig. 11(a), we see that the ratio $\delta_{\rm v}/\Delta_{\rm b}$ is equal to $1/2$ for the two critical values $\kappa_{c_{1}}$ and $\kappa_{c_{2}}$, represented by the red circle and triangle, respectively. In Table 1, we see that the energy gap of the vortex-free ($\rho=0$) background (roughly equal to $\Delta_{\rm b}$) vanishes as $\kappa=0$ and is of order $1$ as $\kappa\to\infty$. Thus the low-energy vortex bands are only well-separated from the background as $\kappa_{c_{1}}<\kappa<\kappa_{c_{2}}$. For less dense triangular vortex configurations in the $\rho\to 0$ limit, we find that this region of validity increases, such that $\kappa_{c_{1}}\to 0$ and $\kappa_{c_{2}}\to\infty$.
@@ B.1 One contact point
This type of transition occurs when the vortex Chern number changes from $\nu_{\rm v}=-1$ to $\nu_{\rm v}=0$ for the dual odd vortex configurations of density $\rho=(n-1)/n$ such that $n$ is not a multiple of three. That is, for densities such as $\rho=12/13$, $18/19$, $24/25$, and etc. These transitions generally occur when $\kappa_{c}\sim 0.1-0.2$. Before the transition ($\kappa<\kappa_{c}$), the system features two gapped Dirac points of opposite chirality and is a Chern insulator with $\nu_{\rm v}=-1$. As $\kappa$ increases, the Dirac points move and at a critical $\kappa_{c}$ they meet at a time-reversal invariant momentum (TRIM) where they annihilate and the gap closes. Exactly at the transition, there is a single contact point that does not carry a topological charge. After this merging transition, the gap reopens and the system becomes a trivial insulator with $\nu_{\rm v}=0$. This merging transition is similar to the one occurring in the Kitaev honeycomb model when evolving from the isotropic point $J_{x}=J_{y}=J_{z}$ to a very anisotropic case [5], or also in uniaxially strained graphene [34].
@@
We note that this Hamiltonian results from a different gauge choice for the $s_{jk}^{(\alpha)}$, $s_{jk}^{(\perp)}$, and $\widetilde{s}_{jk}^{(\perp)}$ variables than what is shown in Fig. 8. In Fig. 12, we plot the dispersion of both this effective model and the low-lying vortex bands of the Kitaev honeycomb model of the $\nu_{\rm v}=-1\to 0$ transition in the $\rho=12/13$ dual odd vortex configuration as $\kappa_{c}=0.219461$ [17]. The proper parameter values for the effective model are given in the caption. The critical point of the effective model appears when $t_{\perp 1}=\epsilon/6$. As $t_{\perp 1}<\epsilon/6$ the Chern number of the effective model is $\nu_{\rm v}=-1$, and when $t_{\perp 1}>\epsilon/6$ the Chern number is $\nu_{\rm v}=0$. The sign of $t_{\perp 2}$ controls the sign of $\nu_{\rm v}$, and the choice of signs of the cosine terms of $\widetilde{g}_{1}(\bm{k})$ controls the position of the contact point within the Brillouin zone. If we were to instead analyze a different vortex density exhibiting a one contact point transition, one may have to alter the signs of the cosine terms in order to properly match the location of the contact point within the Brillouin zone (this is true for the two and six contact points models as well). Note that because the value of $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.38$ is relatively high (i.e., the vortex bands are not so well separated from the background bands), the bandwidth of the effective model overestimates that of the low-lying vortex bands of the Kitaev honeycomb model.
@@ B.2 Two contact points
In Fig. 13, we plot the dispersion of both this effective model and the low-lying vortex bands of the Kitaev honeycomb model of the $\nu_{\rm v}=-2\to 0$ transition in the $\rho=20/21$ dual odd vortex configuration, as $\kappa_{c}=0.106486$. The proper parameter values for the effective model are given in the caption. The critical points of the effective model appear when $t_{\perp 1}=-\epsilon/3$. As $t_{\perp 1}>-\epsilon/3$ the Chern number of the effective model is $\nu_{\rm v}=0$, and when $t_{\perp 1}<-\epsilon/3$ the Chern number is $\nu_{\rm v}=-2$. The sign of $t_{2}$ controls the sign of $\nu_{\rm v}$. Note that $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.62$, and hence the bandwidth of the effective model is slightly larger than that of the low-lying vortex bands of the Kitaev honeycomb model.
We note that in this system, the energy spectrum is gapped and topologically nontrivial only if the sign of one of the sine terms in $\widetilde{g}_{2}(\bm{k})$ is dissimilar to the sign of the other two. If every sine term were positive, for example, then the $2\times 2$ matrix Hamiltonian may be expressed as the linear combination of only two Pauli matrices and would either be gapless before the transition (i.e., a topological semimetal with gapless Dirac points) or gapped after the transition with a trivial Chern number of $\nu_{\rm v}=0$, due to a merging transition.
@@ B.3 Six contact points
In Fig. 14, we plot the dispersion of both this effective model and the low-lying vortex bands of the Kitaev honeycomb model of the $\nu_{\rm v}=-3\to 3$ transition in the $\rho=6/7$ dual odd vortex configuration, as $\kappa_{c}=1.80057$. The proper parameter values for the effective model are given in the caption. The critical points of the effective model appear when $t_{\perp 2}=0$. As $t_{\perp 2}>0$ the Chern number of the effective model is $\nu_{\rm v}=-3$, and when $t_{\perp 2}<0$ the Chern number is $\nu_{\rm v}=3$. The ratio $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.38$ indicates that the bandwidth of the effective model is slightly larger than that of the low-lying vortex bands of the Kitaev honeycomb model. Note that because $t_{1}\neq t_{2}=0$, we do not have the same issue of the sine’s signs as discussed in the two contact points section, and the $2\times 2$ matrix Hamiltonian is still composed of three different Pauli matrices.